# Turning the Tide on Dark Pools? Towards Multi-Stakeholder Vulnerability Notifications in the Ad-Tech Supply Chain

### v1

@@ Abstract
Online advertising relies on a complex and opaque supply chain that involves multiple stakeholders, including advertisers, publishers, and ad-networks, each with distinct and sometimes conflicting incentives. Recent research has demonstrated the existence of ad-tech supply chain vulnerabilities such as dark pooling, where low-quality publishers bundle their ad inventory with higher-quality ones to mislead advertisers. We investigate the effectiveness of vulnerability notification campaigns aimed at mitigating dark pooling. Prior research on vulnerability notifications has primarily focused on single-stakeholder scenarios, and it is unclear whether vulnerability notifications can be effective in the multi-stakeholder ad-tech supply chain. We implement an automated vulnerability notification pipeline to systematically evaluate the responsiveness of various stakeholders, including publishers, ad-networks, and advertisers to vulnerability notifications by academics and activists. Our nine-month long multi-stakeholder notification study shows that notifications are an effective method for reducing dark pooling vulnerabilities in the online advertising ecosystem, especially when targeted towards ad-networks. Further, the sender reputation does not impact responses to notifications from activists and academics in a statistically different way. In addition to being the first notification study targeting the online advertising ecosystem, we are also the first to study multi-stakeholder context in vulnerability notifications.
@@ 1 Introduction
Fraud is rampant in the online advertising supply chain. In 2023, marketers globally spent more than 300 billion dollars on online advertising [1]. But nearly a quarter of this ad spend is lost to ad fraud [2]. Ad fraud generally takes one of three forms: inorganic or fraudulent interactions with ad content [3, 4, 5], manipulating conversion attribution [6, 7], and ad inventory laundering [8, 9, 10, 11]. While the vulnerabilities exploited by each form of ad fraud are multifaceted, they generally exploit the complexity and opacity of the online advertising supply chain [12]. This paper focuses on an emerging type of ad inventory laundering called dark pooling. At a high-level, dark pooling allows low-quality (i.e., brand-unsafe) publishers make their ad inventory indistinguishable, to advertisers or brands, from the ad inventory of higher-quality publishers. Recent work [8] has shown that dark pooling is widespread and helps fund low-quality websites known for publishing misinformation and other brand-unsafe content.
@@
Current strategies for mitigating dark pooling are ineffective. Solutions to prevent the laundering of low-quality ad inventory facilitated by dark pooling generally take the form of new standards for improving transparency and facilitating inventory verification in the online advertising supply chain. For instance, the Interactive Advertising Bureau (IAB) introduced standards such as ads.txt [13], sellers.json [14], ads.cert [15], and the Supply Chain Object [16] to help mitigate this type of ad fraud. Unfortunately, prior research consistently shows that entities do not comply with these standards [17, 18, 8, 19, 20]. Consequently, these measures fail to mitigate, or even detect, ad inventory fraud. Further, given the largely self-regulated nature of the ad-tech ecosystem, we cannot expect significant improvements in compliance rates [21, 22, 23]. Therefore, it is important to explore alternative approaches to mitigate ad inventory fraud. This paper examines the effectiveness of a different tactic, previously unused in the ad-tech supply chain: a vulnerability notification campaign. Specifically, we examine the effectiveness of notification campaigns in mitigating dark pooling.
Findings from prior vulnerability notification research are unlikely to carry over to the online advertising supply chain. Running vulnerability notification campaigns to address security vulnerabilities is not a new idea. In the context of the Web, researchers have conducted many studies on how different notification campaign strategies affect the resolution of various vulnerabilities observed in online services [24, 25, 26, 27, 28, 29, 30, 31]. Generally, these studies show that well-designed notifications can effectively resolve online vulnerabilities when successfully delivered to their targets. These prior studies typically focus on single-stakeholder scenarios — i.e., researchers sent vulnerability notifications to an entity that is capable of resolving the issue (e.g., a website operator) and monitored that entity’s actions. However, this approach and its findings may not apply to the more complex multi-stakeholder ad-tech supply chain. In the ad-tech multi-stakeholder context: (1) vulnerability resolution may require cooperation from multiple independent entities (e.g., publishers, ad-networks, and advertisers); (2) it is unlikely for any one entity to have complete visibility into the supply chain; and (3) incentives of the independent entities may be different and at odds with each. These differences between single- and multi-stakeholder scenarios necessitate evaluating the effectiveness of notification campaigns for the latter. Our work fills this gap by applying (and evaluating) lessons learned from the single-stakeholder context to the multi-stakeholder ad-tech supply chain context.
@@
Developing a vulnerability notification campaign for the ad-tech supply chain is non-trivial. Our research addresses several challenges related to the opacity of the ad-tech supply chain and the multi-stakeholder setting. First, we need to automate the detection of dark pooling vulnerabilities to conduct notifications at a reasonable scale. We address this challenge using methods from prior research [8] to identify dark pooling and the specific entities involved, such as publishers, ad-networks, and advertisers. Second, we must tailor our notifications to make them understandable and useful to each type of stakeholder. We accomplish this by designing notifications that provide entity-specific information, including high-level descriptions of the detected dark pooling vulnerability, technical evidence, and potential remedial actions. Finally, we need to assess the effectiveness of our notifications with statistical rigor and correctly attribute vulnerability resolutions to different stakeholders. We achieve this by conducting our study over a nine-month period and in multiple phases, each focused on inferring the actions of one type of stakeholder.
@@
Our study design and findings advance both notification research and ad fraud research. We first design a multi-phase email-based notification campaign for mitigating dark pooling in the multi-stakeholder ad-tech supply chain (Section 3). This design provides a template for future multi-stakeholder notification studies. Next, we operationalize this design to identify and notify 2.9K entities (1.7K publishers, 644 ad-networks, and 635 advertisers) involved in and affected by dark pooling vulnerabilities. We analyze their responses to answer the following research questions:
@@
  • RQ1. What are the differences in stakeholder attitudes and responses towards dark pooling vulnerability notifications? (Section 4) We conduct a thematic analysis of the stakeholder responses to our notifications. Our approach uncovers three response themes: (1) effort towards resolution; (2) demonstration of concern or awareness; and (3) lack of trust in our data or lack of resources for resolution.
@@
  • RQ2. How does the choice of notification recipient influence the resolution of dark pooling vulnerabilities? (Section 5.2) We measure the rates of dark pooling vulnerability resolution when notifications are sent to publishers, ad-networks, and advertisers. We find that sending notifications to ad-networks has the highest influence on dark pooling vulnerability resolution with 81.6% of notified entities remediating dark pools in their network and 76.9% in partner ad-networks. Publishers demonstrate the least effectiveness to notifications, however, with the highest effect size.
@@
  • RQ3. How does the notification sender influence the resolution of dark pooling vulnerabilities? (Section 5.3) We compare the efficacy of notifications sent from academics with those from activists with a long history of public advocacy in addressing online advertising vulnerabilities [32, 33, 34]. We find that notifications sent as activists are generally as effective as those from academics. The only exception is when the notifications are sent to publishers. In this case, activists have a statistically significantly stronger positive effect on the resolution of dark pools.
@@ 2 Background
In this section, we provide an overview of the online advertising supply chain (Section 2.1), dark pooling vulnerabilities (Section 2.2), and vulnerability notification research (Section 2.3).
@@ 2.1 The online advertising supply chain
How the online advertising supply chain works. Publishers, ad-networks, and advertisers collaborate to create the ad supply chain. When a user visits a publisher’s website, the ad inventory associated with that visit is auctioned off at an ad-network (possibly by other ad-network intermediaries called supply-side platforms). Advertisers then make bids on the auctioned ad inventory available at the ad-network (with the help of brokers called demand-side platforms). Finally, the ad-network places the ad from the winning advertiser (e.g., the highest bid) on the publisher’s website. The advertiser transfers the payment to the ad-network (possibly through intermediary networks), and deposits a portion of the billed impressions into the publisher’s account (possibly through intermediary networks). A simplified version of this process is illustrated in Figure 1.
@@
Challenges associated with the multi-stakeholder supply chain. One consequence of this decentralized multi-stakeholder scenario is that vulnerability resolution is rarely possible without collaboration between multiple, each with their own controls and operating incentives. This interdependence and necessary trust, despite conflicting objectives, between different stakeholders makes the online advertising supply chain a uniquely interesting case-study for assessing the effectiveness of vulnerability notifications.
@@ 2.2 Dark pooling vulnerabilities
How pools become dark? A pool becomes a dark pool when the publishers sharing a single seller ID are organizationally-unrelated and have different reputations [8]. Dark pooling allows low-quality publishers to disguise their inventory as high-quality, leading advertisers to unintentionally purchase low-quality, brand-unsafe inventory. Malicious publishers or ad-networks can facilitate dark pooling. Simple examples of both types of dark pooling vulnerabilities are shown in Figure 2 and explained below. More comprehensive explanations can be found in Vekaria et al. [8] and Papadogiannakis et al. [18]. The remainder of our paper is focused on using notification campaigns to capture the motivations, control, and capabilities of each stakeholder in addressing inventory fraud due to dark pooling.
@@
Dark pooling facilitated by publishers. Recall that publishers aim to join large ad-networks to maximize revenue, which is challenging for low-quality publishers. Dark pooling allows these publishers to misrepresent their existing relationships in order to gain access to large ad-networks. For example, a low-quality publisher (attacker.com) manipulates its ads.txt file to intentionally hide or misrepresent its ad inventory, making it appear as though it belongs to a reputable publisher (victim.com). Next, they change the inventory source to victim.com in future bid requests. An unwitting brand (advertiser.com) might then bid on this inventory, incorrectly believing that it belongs to victim.com (even after using victim.com’s ads.txt file for source verification). Although attacker.com may not immediately financially benefit from the inventory sold by this manipulation (payments go to victim.com’s account), it is useful for demonstrating a higher value for the ad inventory on its website to other ad-networks. Such manipulation is possible because of: (1) misrepresentations, incompleteness, and inaccuracies within ads.txt and sellers.json files made available by publishers and ad-networks, respectively; and (2) poor verification procedures by ad-networks which are incentivized to maximize the sale of ad inventory. Vekaria et al. [8] provide a detailed explanation of these manipuations.
@@
Dark pooling facilitated by ad-networks. Intermediary ad-networks (intermediary.com) provide ‘inventory management’ services to smaller publishers unable to manage their own inventory with in-house teams. Dark pooling occurs when these intermediaries share a single seller ID among clients (i.e., publishers) with different reputations. In this case, ad inventory owned by victim.com and attacker.com auctioned through the intermediary.com’s network appears to an advertiser as owned by intermediary.com and can be made indistinguishable. Ad-networks are not incentivized to prevent dark pooling as it would reduce the volume of inventory traded.
@@
Industry efforts to mitigate dark pooling. The ad-tech ecosystem has introduced several standards to increase supply chain transparency and mitigate inventory fraud. Notable among these are the ads.txt standard [13], sellers.json standard [14], and the RTB SCO (Supply Chain Object) [16] all introduced by the Interactive Advertising Bureau.
@@
The ads.txt standard. Each publisher participating in real-time bidding is expected to implement and maintain an ads.txt file [13] at the root of their domain. This file lists all the ad-network seller domains (and seller IDs assigned to the publisher by the corresponding ad-network) that are authorized to sell or resell the publisher’s inventory.
@@
The sellers.json standard. Ad-networks must maintain a sellers.json file [14] at their domain’s root. This JSON file details all publishers and intermediary ad-networks they work with, along with corresponding domains and ad-network-assigned seller IDs.
@@
Unfortunately, prior work has shown that each of these standards has low adoption rates and serious misrepresentation issues which make them unsuitable for preventing inventory fraud [8, 9, 18, 17, 35].
Related academic research. Recent academic research has focused on measuring the prevalence of inventory fraud. Notably, Kline et al. [10] illustrated several supply chain ad attribution exploits which allowed malicious publishers to inflate their ad revenue by making their inventory appear premium. Papadogiannakis et al. [9, 18] showed how problematic websites exploited supply chain complexities to monetize their website using hidden content, pooling, and identity masquerading. Most closely related to this work, Vekaria et al. [8] provided methods to identify dark pooling, measured its prevalence, and showed how it was used to monetize misinformation publishers.

### v2

# Towards Multi-Stakeholder Vulnerability Notifications in the Ad-Tech Supply Chain

@@ Abstract
Online advertising relies on a complex and opaque supply chain that involves multiple stakeholders, including advertisers, publishers, and ad-networks, each with distinct and sometimes conflicting incentives. Recent research has demonstrated the existence of ad-tech supply chain vulnerabilities such as dark pooling, where low-quality publishers bundle their ad inventory with higher-quality ones to mislead advertisers. We investigate the effectiveness of vulnerability notification campaigns aimed at mitigating dark pooling. Prior research on vulnerability notifications have primarily explored single-stakeholder contexts, leaving multi-stakeholder scenarios understudied. There is limited attention to complex multi-stakeholder supply chain ecosystems such as ad-tech supply chain, where resolving vulnerabilities often requires coordinated action across entities with misaligned incentives and interdependent roles. We address this gap by implementing the first online advertising supply chain vulnerability notification pipeline to systematically evaluate the responsiveness of various stakeholders in ad-tech supply chain, including publishers, ad-networks, and advertisers to vulnerability notifications by academics and activists. Our nine-month long automated multi-stakeholder notification study shows that notifications are an effective method for reducing dark pooling vulnerabilities in the online advertising ecosystem, especially when targeted towards ad-networks. Further, the sender reputation does not impact responses to notifications from activists and academics in a statistically different way. Overall, our research fosters industry-scale solution to combat ad inventory fraud and fosters future research on feasibility of multi-stakeholder vulnerability notifications in other supply chain ecosystems.
@@ 1 Introduction
Fraud is rampant in the online advertising supply chain. In 2023, marketers globally spent more than 300 billion dollars on online advertising [ad-spend-bechmarks-2023]. But nearly a quarter of this ad spend is lost to ad fraud [juniper]. Ad fraud generally takes one of three forms: inorganic or fraudulent interactions with ad content [pearce-ccs14, kshetri-sp10, springborn-security13], manipulating conversion attribution [chachra-imc2015, stone-imc2011], and ad inventory laundering [vekaria2023inventory, papadogiannakis2023funds, kline2022placement, inventorymisrepresentation]. While the vulnerabilities exploited by each form of ad fraud are multifaceted, they generally exploit the complexity and opacity of the online advertising supply chain [supplychaintransparency]. This paper focuses on an emerging type of ad inventory laundering called dark pooling. At a high-level, dark pooling allows low-quality (i.e., brand-unsafe) publishers make their ad inventory indistinguishable, to advertisers or brands, from the ad inventory of higher-quality publishers. Recent work [vekaria2023inventory] has shown that dark pooling is widespread and helps fund low-quality websites known for publishing misinformation and other brand-unsafe content.
@@
Current dark pooling mitigation strategies are ineffective. Solutions to prevent the laundering of low-quality ad inventory facilitated by dark pooling generally take the form of new standards for improving transparency and facilitating inventory verification in the online advertising supply chain. For instance, the Interactive Advertising Bureau (IAB) introduced standards such as ads.txt [adstxt-specification], sellers.json [sellersjson-specification], and the Supply Chain Object [openrtb-guidelines] to help mitigate this type of ad fraud. Unfortunately, prior research consistently shows that entities do not comply with these standards [bashir2019longitudinal, papadogiannakis2025welcome, vekaria2023inventory, tingleff2019three, pastor2020establishing]. Consequently, these measures fail to mitigate, or even detect, ad inventory fraud. Further, given the largely self-regulated nature of the ad-tech ecosystem, we cannot expect significant improvements in compliance rates [zukina2012accountability, culnan2000protecting, hirsch2010law]. Therefore, it is important to explore alternative approaches to mitigate ad inventory fraud. This paper examines the effectiveness of a different tactic, previously unused in the ad-tech supply chain: a vulnerability notification campaign. Specifically, we examine the effectiveness of notification campaigns in mitigating dark pooling.
Findings from prior vulnerability notification research are unlikely to carry over to the online advertising supply chain. In context of the Web, researchers have conducted many studies on how different notification campaign strategies affect the resolution of various vulnerabilities observed in online services [stock2016hey, li2016you, li2016remedying, cetin2016understanding, cetin2017make, ccetin2019tell, maass2021snail, maass2021effective]. These prior works primarily focus on single-stakeholder scenarios, where vulnerability notifications are sent to a single entity that is capable of resolving it (e.g., a website operator). A few studies have examined multi-stakeholder, but typically in contexts where the vulnerability (1) is associated with a single stakeholder (e.g., WHOIS and IP WHOIS notified about non-secure DNS vulnerability associated solely with nameserver operator [cetin2017make]) or (2) does not require coordinated effort from two or more notified stakeholders to fix the vulnerability (e.g., either hosting providers or CERTs can independently perform remediations to handle the notification about vulnerable websites, without a need to coordinate [stock2016hey]) However, there remains limited understanding of the feasibility and effectiveness of vulnerability notifications in complex and opaque multi-stakeholder ecosystems such as ad-tech supply chain, where (1) vulnerability could exist at multiple stakeholders. (2) vulnerability resolution may require cooperation from multiple independent entities (e.g., publishers may need to coordinate with ad-networks to fix the dark pooling vulnerability); (3) incentives of the independent entities may be different and at odds with each other. Designing notification strategies for such ecosystems requires incorporating these interdependencies and incentive misalignments into the study design. Our work fills this gap by studying multi-stakeholder ad-tech supply chain ecosystem for the first time.
@@
Developing a vulnerability notification campaign for the ad-tech supply chain is non-trivial. Our research addresses several challenges related to the opacity of the multi-stakeholder ad-tech supply chain. First, we need to automate the detection of dark pooling vulnerabilities to conduct notifications at a reasonable scale. We address this challenge using methods from prior research [vekaria2023inventory] to identify dark pooling and the specific entities involved, such as publishers, ad-networks, and advertisers. Second, we must tailor our notifications to make them understandable and useful to each type of stakeholder. We accomplish this by designing notifications that provide entity-specific information, including high-level descriptions of the detected dark pooling vulnerability, technical evidence, and potential remedial actions. Finally, we need to assess the effectiveness of our notifications with statistical rigor and correctly attribute vulnerability resolutions to different stakeholders. We achieve this by conducting our study over a nine-month period and in multiple phases, each focused on inferring the actions of one type of stakeholder.
@@
Our study design and findings advance both notification research and ad fraud research. We first design a multi-phase email-based notification campaign for mitigating dark pooling in the multi-stakeholder ad-tech supply chain (Section 3). This design provides a template for future fragmented multi-stakeholder notification studies. Next, we operationalize this design to identify and notify 2.9K entities (1.7K publishers, 644 ad-networks, and 635 advertisers) involved in and affected by dark pooling vulnerabilities. We analyze their responses to answer the following research questions:
@@
$\bullet$ RQ1. What are the differences in stakeholder attitudes and responses towards dark pooling vulnerability notifications? (Section 4) We conduct a thematic analysis of the stakeholder responses to our notifications. Our approach uncovers three response themes: (1) effort towards resolution; (2) demonstration of concern or awareness; and (3) lack of trust in our data or lack of resources for resolution.
@@
$\bullet$ RQ2. How does the choice of notification recipient influence the resolution of dark pooling vulnerabilities? (Section 5.2) We measure the rates of dark pooling vulnerability resolution when notifications are sent to publishers, ad-networks, and advertisers. We find that sending notifications to ad-networks has the highest influence on dark pooling vulnerability resolution with 81.6% of notified entities remediating dark pools in their network and 76.9% in partner ad-networks. Publishers demonstrate the least effectiveness to notifications, however, with the highest effect size.
@@
$\bullet$ RQ3. How does the notification sender influence the resolution of dark pooling vulnerabilities? (Section 5.3) We compare the efficacy of notifications sent from academics with those from activists with a long history of public advocacy in addressing online advertising vulnerabilities [ferraz2021sleeping, braun2019activism, checkmyads]. We find that notifications sent as activists are generally as effective as those from academics. The only exception is when the notifications are sent to publishers. In this case, activists have a statistically significantly stronger positive effect on the resolution of dark pools.
@@ 2.1 The online advertising supply chain
How the online advertising supply chain works. Publishers, ad-networks, and advertisers collaborate to create the ad supply chain. When a user visits a publisher’s website, the ad inventory associated with that visit is auctioned off at an ad-network (possibly by other ad-network intermediaries called supply-side platforms). Advertisers then make bids on the auctioned ad inventory available at the ad-network (with the help of brokers called demand-side platforms). Finally, the ad-network places the ad from the winning advertiser (e.g., the highest bid) on the publisher’s website. The advertiser transfers the payment to the ad-network (possibly through intermediary networks), and deposits a portion of the billed impressions into the publisher’s account (possibly through intermediaries). A simplified version of this process is illustrated in Figure 1.
@@
Challenges with the multi-stakeholder supply chain. One consequence of this decentralized multi-stakeholder scenario is that vulnerabilities can exist at multiple stakeholders and resolution is rarely possible without collaboration between multiple entities, each with their own controls and operating incentives. This interdependence and necessary trust, despite conflicting objectives, between different stakeholders makes the online advertising supply chain a uniquely interesting case-study for assessing the effectiveness of vulnerability notifications.
@@ 2.2 Dark pooling vulnerabilities
How pools become dark? A pool becomes a dark pool when the publishers sharing a single seller ID are organizationally-unrelated and have different reputations [vekaria2023inventory]. Dark pooling allows low-quality publishers to disguise their inventory as high-quality, leading advertisers to unintentionally purchase low-quality, brand-unsafe inventory. Malicious publishers or ad-networks can facilitate dark pooling. Simple examples of both types of dark pooling vulnerabilities are shown in Figure 2 and explained below. More comprehensive explanations can be found in Vekaria et al. [vekaria2023inventory] and Papadogiannakis et al. [papadogiannakis2025welcome]. The remainder of our paper is focused on using notification campaigns to capture the motivations, control, and capabilities of each stakeholder in addressing inventory fraud due to dark pooling.
@@
Dark pooling facilitated by publishers. Recall that publishers aim to join large ad-networks to maximize revenue, which is challenging for low-quality publishers. Dark pooling allows these publishers to misrepresent their existing relationships in order to gain access to large ad-networks. For example, a low-quality publisher (attacker.com) manipulates its ads.txt file to intentionally hide or misrepresent its ad inventory by including entries from ads.txt of some reputable publisher (victim.com), making attacker’s inventory appear as though it belongs to a reputable publisher. Next, they change the inventory source to victim.com in future bid requests. An unwitting brand (advertiser.com) might then bid on this inventory, incorrectly believing that it belongs to victim.com (even after using victim.com’s ads.txt file for source verification). Although attacker.com may not immediately financially benefit from the inventory sold by this manipulation (payments go to victim.com’s account), it is useful for demonstrating a higher value for the ad inventory on its website to other ad-networks. Such manipulation is possible because of: (1) misrepresentations, incompleteness, and inaccuracies within ads.txt and sellers.json files made available by publishers and ad-networks, respectively; and (2) poor verification procedures by ad-networks which are incentivized to maximize the sale of ad inventory. Vekaria et al. [vekaria2023inventory] discuss these manipuations in detail.
@@
Dark pooling facilitated by ad-networks. Intermediary ad-networks (intermediary.com) provide ‘inventory management’ services to smaller publishers unable to manage their own inventory with in-house teams. Dark pooling occurs when these intermediaries share a single seller ID among clients (i.e., publishers) with different reputations. In this case, ad inventory owned by victim.com and attacker.com auctioned through the intermediary.com’s network appears to an advertiser as owned by intermediary.com and can be made indistinguishable (for e.g., attacker.com misrepresents its inventory source as victim.com). Ad-networks are not incentivized to prevent dark pooling as it would reduce the volume of inventory traded.
In summary, our dark pooling threat model may constitute following scenarios:
❶ Ad-network bundles ad inventory from different publishers and sells them to buyers using a common sellerID owned by the ad-network. Given the challenges with source verification, ad-network (knowingly or unknowingly) sells misrepresented inventory (i.e., a problematic publisher (mis)represented as victim[.]com).
❷ Problematic and victim publishers may (not) collude:
$\bullet$ If they do not collude: Problematic publishers may still improve their reputation by using ad-network assigned victim publishers’ IDs (e.g., to show that they work with reputable ad-networks to convince more advertisers to bid or gain access to new ad-networks) as a result of pooling.
$\bullet$ If they collude: Problematic and victim publishers could have some sort of a revenue sharing model where problematic publishers share revenue generated by using victim’s seller ID with the victim publisher. However, if an advertiser detects low ROI or ad-misplacements associated with certain pooled seller ID, they may block it – hurting victim publisher’s long term revenue.
@@
Industry efforts to mitigate dark pooling. The ad-tech ecosystem has introduced several standards to increase supply chain transparency and mitigate inventory fraud. Notable among these are the ads.txt standard [adstxt-specification], sellers.json standard [sellersjson-specification], and the RTB (real-time bidding) SCO (Supply Chain Object) [openrtb-guidelines] all introduced by the Interactive Advertising Bureau.
@@
The ads.txt standard. Each publisher participating in RTB is expected to implement and maintain an ads.txt file [adstxt-specification] at the root of their domain. This file lists all ad-network seller domains (and seller IDs assigned to the publisher by the corresponding ad-network) that are authorized to sell or resell the publisher’s inventory.
@@
The sellers.json standard. Ad-networks must maintain a JSON file – sellers.json [sellersjson-specification] at their domain’s root. This file details all publishers and intermediary ad-networks they work with, along with corresponding domains and ad-network-assigned seller IDs.
@@
Unfortunately, prior work has shown that these standards have relatively low adoption rates and serious misrepresentation issues [vekaria2023inventory, papadogiannakis2023funds, papadogiannakis2025welcome, bashir2019longitudinal, olejnik2018enhancing]. Moreover, lack of enforcement makes them unsuitable for preventing inventory fraud by itself.
Related academic research. Recent academic research has focused on measuring the prevalence of inventory fraud. Notably, Kline et al. [kline2022placement] illustrated several supply chain ad attribution exploits which allowed malicious publishers to inflate their ad revenue by making their inventory appear premium. Papadogiannakis et al. [papadogiannakis2023funds, papadogiannakis2025welcome, papadogiannakis2024ad] showed how problematic websites exploited supply chain complexities to monetize their website using hidden content, pooling, and identity masquerading. Most closely related to this work, Vekaria et al. [vekaria2023inventory] provided methods to identify dark pooling, measured its prevalence, and showed how it monetizes misinformation publishers.